# Soft-CAM: Making black box models self-explainable for high-stakes decisions

### v1

@@ D.2 Activation consistency
We quantify activation consistency only for the SoftCAM variants, as post-hoc methods are not inherently explainable, meaning their explanations do not directly influence the model’s decision-making process.
@@
The results align well with qualitative visualizations. On the Fundus dataset, the sparse SoftCAM model exhibits a higher proportion of positive activations with the ResNet backbone, attributed to reduced false positives from the dense model, and fewer negative activations, reflecting the suppression of low-importance activations to zero. On the VGG backbone, regularization primarily reduces false-positive activations from the dense model but leads to a slight increase in activations on healthy samples. Similar result can be observed on the RSNA dataset.
@@
On the OCT dataset, the dense SoftCAM with the ResNet backbone generally produces coarse-grained evidence around lesion areas. In contrast, the sparse variant refines these explanations, resulting in lower positive and negative activations across both disease and healthy samples, suggesting more selective and focused localization. However, with the VGG backbone, a higher proportion of negative activations is observed, reflecting the impact of the regularization strength, highlighting the importance of appropriately tuning this parameter for different architectures.
@@
D.3 Precision and sensitivity analysis
We quantitatively evaluate the explanations generated by various methods using the ResNet and VGG backbones on the RSNA dataset. With the ResNet model, the dense SoftCAM achieves the highest localization precision, whereas the sparse SoftCAM yields the best results in terms of sensitivity. This discrepancy underscores the importance of developing evaluation metrics that balance human-aligned localization quality with model fidelity, capturing both interpretability and decision relevance.
@@ D.3 Precision and sensitivity analysis
D.4 SoftCAM provides localized and faithful explanations
Table 4: Top-k localization precision and sensitivity. Sensitivity is quantified as the Area Under the Deleted Curve (AUDC), where lower values indicate greater faithfulness—that is, a larger drop in the model’s confidence when the most relevant patches are removed. For precision, higher values indicate better alignment between saliency maps and ground truth annotations. We refer to AUDC as “Del” and Top-K as “Top” with $K=30$.
ResNet (Topk $\uparrow$, AUDC $\downarrow$) VGG (Topk $\uparrow$, AUDC $\downarrow$)
@@ D.4 SoftCAM provides localized and faithful explanations
Itgd Grad. $0.28$ $0.63$ $0.20$ $0.70$ $0.52$ $0.98$ $0.33$ $\bf{0.62}$ $0.2$ $0.51$ $0.55$ $0.88$
dense SoftCAM $0.39$ $0.69$ $0.46$ $0.61$ $0.65$ $\bf{0.92}$ $0.54$ $0.63$ $0.72$ $0.45$ $0.64$ $0.84$
@@
Appendix E Activation precision and sensitivity on the RSNA dataset
E.1 Lasso vs Ridge penalty
Figure 11: Model selection on validation sets under varying Lasso and Ridge regularization strengths. The regularization coefficients $\lambda_{1}$ and $\lambda_{2}$ influence model performance. The red markers indicate the selected regularization values, chosen to balance classification performance.
E.2 Activation precision vs. activation sensitivity
Table 5: Activation Precision (AP) vs. Activation Sensitivity (AS) for different SoftCAM variants and baseline post-hoc methods. The dense SoftCAM consistently lies between the lasso and ridge variants, highlighting the importance of balancing the two regularization terms to achieve an optimal trade-off between precision and completeness in the explanations.
@@ E.1 Lasso vs Ridge penalty
lasso SoftCAM $0.440$ $\bf{0.316}$ $0.412$ $\bf{0.396}$
E.3 More examples: activation precision vs activation sensitivity
Figure 12: Additional examples of localization evaluation on the RSNA dataset for pneumonia detection. Each column shows explanation maps generated by different methods. Ground-truth bounding boxes are overlaid on each map, with the top-right value indicating the activation precision, while the top-left value indicates the activation sensitivity. The high precision from the lasso model and the more complete explanations from the ridge model emphasize the importance of balancing the two regularization terms to achieve an optimal trade-off.
Appendix F Multi-class analysis
F.1 Regularization
@@
Figure 13: Model selection on validation sets under varying Lasso regularization strengths. The regularization coefficients $\lambda_{1}$ influence model performance. The red markers indicate the selected regularization values to balance classification performance.
F.2 Faithfulness
As no ground-truth lesion annotations were available for the multi-class tasks, we evaluated the faithfulness of the explanations by measuring their contribution to model predictions. For correctly classified test samples, we progressively removed top-k ($k=30$) ranked patches (based on the explanation maps) and tracked the average drop in class confidence (Fig. 14).
Figure 14: Sensitivity analysis.
F.3 Area Under the Deleted Curve
The area under the deletion curve (AUDC) was computed from the sensitivity analysis (Fig. 14). In both tasks, the dense and sparse SoftCAM achieved superior performance, with sparse SoftCAM yielding the lowest AUDC, indicating the highest faithfulness (Tab. 6).
@@ F.3 Area Under the Deleted Curve
F.4 Qualitative explanation on retinal fundus images
For the multi-class tasks on DR detection from fundus images, SoftCAM variants produced more focused and class-consistent explanations. In addition to the sparse and dense evidence maps, we also provide visualizations for post-hoc methods.
@@ F.4 Qualitative explanation on retinal fundus images
Figure 16: Class-specific explanation with the VGG backbone. The application of our method to multi-class DR detection demonstrates the utility of class-specific explanations produced by the sparse SoftCAM, which more precisely highlight disease-relevant regions compared to the dense SoftCAM and the best-performing post-hoc method, ScoreCAM. In the example shown, the image is labeled as severe DR, and the highlighted regions correspond to suspicious areas, reflecting relevant DR lesions.
F.5 Qualitative explanation on retinal OCT images
For the multi-class tasks on retinal disease classification from OCT images, SoftCAM variants produced more focused and class-consistent explanations. In addition to the sparse and dense evidence maps, we also provide visualizations for GradCAM and Guided BP, as these were the best-performing post-hoc methods for the ResNet and VGG backbones, in terms of the Area Under the Deletion Curve (Tab. 6).
Figure 17: Class-specific explanation with the ResNet backbone. SoftCAM applied to multi-class retinal disease classification demonstrates the utility of class-specific explanations, with the sparse SoftCAM variant more precisely highlighting disease-relevant regions compared to both the dense SoftCAM and the best-performing post-hoc methods, GradCAM and Guided Backpropagation. In the example shown, the image is labeled as Diabetic Macular Edema (DME), and the highlighted regions produced by sparse SoftCAM highlight suspicious areas, reflecting relevant retinal lesions.
Figure 18: Class-specific explanation with the VGG backbone. SoftCAM applied to multi-class retinal disease classification demonstrates the utility of class-specific explanations, with the sparse SoftCAM variant more precisely highlighting disease-relevant regions compared to both the dense SoftCAM and the best-performing post-hoc methods, GradCAM and Guided Backpropagation. In the example shown, the image is labeled as Diabetic Macular Edema (DME), and the highlighted regions produced by sparse SoftCAM highlight suspicious areas, reflecting relevant retinal lesions.

### v2

@@ G.2 Activation consistency
We quantified activation consistency only for the SoftCAM variants, as post-hoc methods are not inherently explainable, meaning their explanations do not directly influence the model’s decision-making process.
@@
The results align well with qualitative visualizations. On the Fundus dataset, the sparse SoftCAM model exhibits a higher proportion of positive activations with the ResNet backbone, attributed to reduced false positives from the dense model, and fewer negative activations, reflecting the suppression of low-importance activations to zero. On the VGG backbone, regularization primarily reduces false-positive activations from the unregularized model but leads to a slight increase in activations on healthy samples. A similar result can be observed on the RSNA dataset.
@@
On the OCT dataset, the unregularized SoftCAM with the ResNet backbone generally produces coarse-grained evidence around lesion areas. In contrast, the sparse variant refines these explanations, resulting in lower positive and negative activations across both disease and healthy samples, suggesting more selective and focused localization. However, with the VGG backbone, a higher proportion of negative activations is observed, reflecting the impact of the regularization strength—highlighting the importance of appropriately tuning this parameter for different architectures.
@@
G.3 Precision and sensitivity analysis
We quantitatively evaluate the explanations generated by various methods using the ResNet and VGG backbones on the RSNA dataset. With the ResNet model, the unregularized SoftCAM achieves the highest localization precision, whereas the sparse SoftCAM yields the best results in terms of sensitivity. This discrepancy underscores the importance of developing evaluation metrics that balance human-aligned localization quality with model fidelity, capturing both interpretability and decision relevance.
@@ G.3 Precision and sensitivity analysis
G.4 SoftCAM provides localized and faithful explanations
Alongside the visualization comparing SoftCAM variants with post-hoc methods, we provide the corresponding quantitative metrics, reporting both top-k precision and faithfulness.
Table 4: Top-15 localization precision and sensitivity. Sensitivity is quantified as the Area Under the Deleted Curve (AUDC), where lower values indicate greater faithfulness. For precision, higher values indicate better alignment between saliency maps and ground truth annotations. We refer to AUDC as “Del” and Top-K as “Top”.
ResNet (Top $\uparrow$, Del $\downarrow$) VGG (Top $\uparrow$, Del $\downarrow$)
Fundus OCT RSNA Fundus OCT RSNA
Top Del Top Del Top Del Top Del Top Del Top Del
ScoreCAM $0.22$ $0.77$ $0.12$ $0.84$ $0.72$ $0.98$ $0.30$ $0.78$ $0.12$ $0.76$ $\bf{0.74}$ $0.93$
LayerCAM $0.25$ $0.76$ $0.11$ $0.85$ $0.75$ $0.97$ $0.30$ $0.76$ $0.11$ $0.76$ $0.72$ $0.91$
GradCAM $0.37$ $0.75$ $0.15$ $0.84$ $0.75$ $0.97$ $\bf{0.65}$ $0.79$ $0.15$ $0.77$ $0.72$ $0.92$
Guided BP $0.38$ $\bf{0.69}$ $0.36$ $0.80$ $0.60$ $0.98$ $0.48$ $\bf{0.72}$ $0.36$ $\bf{0.65}$ $0.66$ $0.92$
Itgd Grad $0.34$ $0.73$ $0.30$ $0.82$ $0.56$ $0.98$ $0.40$ $0.75$ $0.30$ $0.77$ $0.61$ $0.93$
SoftCAM $0.40$ $0.79$ $0.46$ $0.84$ $0.74$ $\bf{0.95}$ $0.54$ $0.73$ $0.72$ $0.68$ $\bf{0.74}$ $0.92$
sparse SoftCAM $\bf{0.52}$ $0.78$ $\bf{0.86}$ $\bf{0.57}$ $\bf{0.78}$ $\bf{0.95}$ $\bf{0.65}$ $0.78$ $\bf{0.82}$ $0.66$ $0.71$ $\bf{0.90}$
Table 5: Top-k localization precision and sensitivity, $k=30$. Sensitivity is quantified as the Area Under the Deleted Curve (AUDC), where lower values indicate greater faithfulness—that is, a larger drop in the model’s confidence when the most relevant patches are removed. For precision, higher values indicate better alignment between saliency maps and ground truth annotations.
ResNet (Top $\uparrow$, Del $\downarrow$) VGG (Top $\uparrow$, Del $\downarrow$)
@@ G.4 SoftCAM provides localized and faithful explanations
Itgd Grad $0.28$ $0.63$ $0.20$ $0.70$ $0.52$ $0.98$ $0.33$ $\bf{0.62}$ $0.2$ $0.51$ $0.55$ $0.88$
SoftCAM $0.39$ $0.69$ $0.46$ $0.61$ $0.65$ $\bf{0.92}$ $0.54$ $0.63$ $0.72$ $0.45$ $0.64$ $0.84$
@@
Appendix H Activation precision and sensitivity on the RSNA dataset
H.1 Lasso vs Ridge penalty
Figure 10: Model selection on validation sets under varying values. The regularization strengths $\lambda_{1}$ and $\lambda_{2}$ influence model performance. The red markers indicate the selected regularization values, chosen to balance classification performance.
H.2 Activation precision vs. activation sensitivity
Table 6: Activation Precision (AP) vs. Activation Sensitivity (AS) for different SoftCAM variants and baseline post-hoc methods. The base SoftCAM consistently lies between the lasso and ridge variants, highlighting the importance of balancing $\ell_{1}$ and $\ell_{2}$ values to achieve an optimal trade-off between precision and completeness.
@@ H.1 Lasso vs Ridge penalty
H.3 More examples: activation precision vs activation sensitivity
Figure 11: Localization precision and sensitivity for pneumonia detection. Each column shows explanations generated by different methods. Ground-truth bounding boxes are drawn on each map, with the top-right value indicating the activation precision, while the top-left value indicates the activation sensitivity. This shows the trade-off between ridge and lasso regularization.
Appendix I Multi-class analysis
We extended our method to the multi-class setting for retinal disease diagnosis, using the same training setup as in the binary tasks. The only modification was adjusting the number of output classes in the evidence layer to 5 for DR grading (fundus dataset) and 4 for retinal disease classification (OCT dataset, and selecting appropriate lasso penalties (e.g. $\lambda_{1}=9.10^{-4}$ vs. $\lambda_{1}=3.10^{-6}$ for ResNet and VGG on the OCT dataset).
I.1 Regularization
@@
Figure 12: Model selection on validation sets under varying sparsity values. The regularization coefficients $\lambda_{1}$ influence model performance. The red markers indicate the selected regularization values to balance classification performance.
I.2 Area Under the Deleted Curve
The relative area under the deletion curve (AUDC) was computed from the sensitivity analysis by occluding the top-30 patches ranked by importance in the explanation map (Fig. LABEL:fig5:multi-class-sensitivity). In both tasks, the dense and sparse SoftCAM achieved superior performance, with sparse SoftCAM yielding the lowest AUDC, indicating the highest faithfulness (Tab. 7).
@@ I.2 Area Under the Deleted Curve
I.3 Qualitative explanation on retinal fundus images
For the multi-class DR detection tasks on fundus images, SoftCAM variants produced more focused and class-consistent explanations. Alongside the sparse and unregularized evidence maps, we also include visualizations from post-hoc methods.
@@ I.3 Qualitative explanation on retinal fundus images
Figure 14: Class-specific explanation with the VGG backbone.
I.4 Qualitative explanation on retinal OCT images
For multi-class OCT-based retinal disease classification, SoftCAM variants yielded more focused and class-consistent explanations. We also provide visualizations from other methods. Note that GradCAM and Guided BP were the best-performing post-hoc methods for ResNet and VGG according to the Area Under the Deletion Curve (Tab. 7).
Figure 15: Class-specific explanation with the ResNet backbone. Sparse SoftCAM provides more precise class-specific localization than unregularized SoftCAM and leading post-hoc methods, highlighting clinically relevant regions.
Figure 16: Class-specific explanation with the VGG backbone.